OpenStudy (bbb911):

The measures of the angels of a triangle are x, 2x, and 3x. What is the measure of the largest angle?

OpenStudy (anonymous):

120

OpenStudy (anonymous):

its 90

OpenStudy (anonymous):

JK 30,6090

OpenStudy (bbb911):

thanks Kanwar

OpenStudy (anonymous):

x+3x+2x=180 6x=180 x=180/6 x=30 Largest angle=3x=3*30=90 which implies a right angled triangle

OpenStudy (bbb911):

thanks

OpenStudy (anonymous):

i need to take a shower and wash all the salt off before i try giving people my rrunners high answers